OpenStudy (anonymous):

Find the indicated sum for the arithmetic series. S16 for 12 + 7 + 2 + (-3) + ...

OpenStudy (ash2326):

a1=12 d=common difference=-5 a16=a1+(16-1)*d a16=12+15*-5 a16=-63 sum of n terms=n(a1+an)/2 n=16 a1=12 a16=-63 sum=16(12-63)/2 sum=8*(-51)=-408

OpenStudy (ash2326):

a1=first term a16=sixteenth term of the series